An aqueous solution of a platinum salt is electrolyzed at a current of 2.498 A for 3.00 h. As a result, 9.09 g of metallic Pt are formed at the cathode. Following this half reaction: Ptn+ + ne- → Pt, calculate the charge (represented by “n+” on the half reaction) on the Pt ions in this solution.

1
Expert's answer
2021-12-27T08:14:02-0500

Quantity of electricity "Q=current (I)\u00d7time(t)"

"Q=2.498×3×3600\\=26978.4C"

1 Faraday"=96485C"

Number of Faradays"=\frac{26978.4}{96485}=0.279612"


Molar mass of Platinum"=195.084g"


"n" Faradays discharges "195.084g \>of \>Pt^{n+} \>"


But "0.279612 \>Faradays\> discharges\> 9.09g"


"\therefore\>\frac{n}{0.279612}=\frac{195.084}{9.09}"


"n=6.00086"

Charge "(n+)=+6"